# An Expanded Massive Multilingual Dataset for High-Performance Language Technologies

### v2

@@ Abstract
Training state-of-the-art large language models requires vast amounts of clean and diverse textual data. However, building suitable multilingual datasets remains a challenge. In this work, we present HPLT v2, a collection of high-quality multilingual monolingual and parallel corpora. The monolingual portion of the data contains 8T tokens covering 193 languages, while the parallel data contains 380M sentence pairs covering 51 languages. We document the entire data pipeline and release the code to reproduce it. We provide extensive analysis of the quality and characteristics of our data. Finally, we evaluate the performance of language models and machine translation systems trained on HPLT v2, demonstrating its value.
@@
High-Performance Language Technologies
Laurie Burchell1, Ona de Gibert2, Nikolay Arefyev3, Mikko Aulamo2, Marta Bañón4, Pinzhen Chen1, Mariia Fedorova3, Liane Guillou1, Barry Haddow1, Jan Hajič5, Jindřich Helcl5, Erik Henriksson6, Mateusz Klimaszewski1, Ville Komulainen6, Andrey Kutuzov3, Joona Kytöniemi6, Veronika Laippala6, Petter Mæhlum3, Bhavitvya Malik1, Farrokh Mehryary6, Vladislav Mikhailov3, Nikita Moghe1, Amanda Myntti6, Dayyán O’Brien1, Stephan Oepen3, Proyag Pal1, Jousia Piha6, Sampo Pyysalo6, Gema Ramírez-Sánchez4, David Samuel3, Pavel Stepachev1, Jörg Tiedemann2, Dušan Variš5, Tereza Vojtěchová5, Jaume Zaragoza-Bernabeu4 1University of Edinburgh, 2University of Helsinki, 3University of Oslo, 4Prompsit Language Engineering, 5Charles University, 6University of Turku Contact: https://hplt-project.org
@@ 1 Introduction
In this paper, we introduce a new set of text corpora extracted from 4.5PB of the Internet Archive (IA) [^1] https://archive.org and Common Crawl (CC) [^2] https://commoncrawl.org, dubbed HPLT v2. We build on the work of de Gibert et al. (2024) (hereafter referred to as HPLT v1.2) with an improved extraction pipeline and a much larger set of input crawls to produce the HPLT v2 collection of monolingual and parallel corpora. To our knowledge, our new corpus is the only large-scale text collection extracted from the IA, apart from HPLT v1.2. We release HPLT v2 under the permissive Creative Commons Zero (CC0) license [^3] https://creativecommons.org/share-your-work/public-domain/cc0/. We do not claim ownership of any of the text from which this data has been extracted. and provide the code to replicate our pipeline. Our main contributions can be summarised as:
  • We compile monolingual corpora covering 193 languages and containing approximately 52 trillion characters and 8 trillion tokens.
@@
Figure 1: The distribution of documents in the HPLT v2 cleaned dataset by language family and language variety. Shortened ISO 639-3 language codes are used for reasons of space.
@@ 2 Related work
The increasing data demands of state-of-the-art large language models have driven a rapid growth in both the number and the size of text corpora. We provide a summary of some well-known collections in Appendix A. Whilst LLMs trained on ostensibly English data have shown impressive multilingual capabilities Armengol-Estapé et al. (2022), of particular relevance to this work is the growing shift towards explicitly multilingual corpora. Compared with earlier efforts (e.g. OSCAR Suárez et al. (2019), CC-100 Conneau et al. (2020a) and mC4 Xue et al. (2021)), more recent multilingual datasets cover increasing numbers of languages, e.g. CulturaX Nguyen et al. (2024a) and MADLAD-400 Kudugunta et al. (2024). HPLT v2 continues this trend by aiming for significant coverage of a wide range of languages. We note that the majority of previous multilingual datasets are sourced from CC, whereas much of HPLT v2 is composed of IA crawls. This means that HPLT v2 can be used in conjunction with these existing datasets as a complementary source.
@@
The closest to the current work is the HPLT v1.2 dataset introduced in de Gibert et al. (2024). Compared to their work, we process more data (21 billion vs. 5 billion documents) using an improved pipeline (Section 4), resulting in a significantly larger dataset (52 trillion characters compared to 42 trillion). HPLT v2 also covers 193 languages compared to 75 in HPLT v1.2. Finally, the HPLT v2 collections are of higher quality than those in HPLT v1.2, as shown through comparative analysis and evaluation (Sections 5 and 6).
@@ 3.1 Monolingual datasets
The monolingual portion of HPLT v2 covers 193 language varieties [^5] Language varieties are labelled with an ISO 639-3 code denoting the variety plus an ISO 15924 four-letter code denoting the script, separated by an underscore: e.g., gla_Latn. and is published in two variants: ‘deduplicated’ (21 TB) and ‘cleaned’ (15 TB). In the latter variant, the documents filtered by our cleaning heuristics (see Section 4.2) are excluded. For training LLMs, we recommend using the cleaned variant, but we also publish the datasets before cleaning (‘deduplicated’) so that it is possible to apply custom cleaning pipelines to the HPLT v2 data. In total, the deduplicated monolingual HPLT v2 datasets contain approximately 7.6 trillion white-space separated tokens and 52 trillion characters, extracted from 21 billion documents. HPLT v2 is published in the JSONL format, with one document per line.
@@ 3.2 Parallel datasets
We use the monolingual HPLT v2 datasets to extract parallel data covering 50 languages paired with English. We aimed for a diverse range of language varieties and scripts in the low to medium resource range (listed in Table 4). We align these to English since this configuration has the highest potential for finding high-quality parallel data. We release our data in both XML and bitext format.
@@
We leverage the English-centric HPLT v2 parallel resources to create a multi-way parallel corpus, obtained by pivoting via English. This corpus includes 1275 language pairs and contains over 16.7 billion parallel sentences.
@@ 4.1 Text extraction from web crawls
This stage of the pipeline extracts the main textual content from HTML pages and groups it into language-specific subsets. It first parses the HTML pages into a tree representation. Next, it removes likely machine-translated texts by searching for indicative HTML tags and attributes. It then removes boilerplate (i.e. parts of a web page that do not contribute to its main content) using Trafilatura 1.8.0 (Barbaresi, 2021). Following hyperparameter experimentation, we set include_comments=False, include_tables=False, no_fallback=False and MIN_EXTRACTED_SIZE=0, with all other hyperparameters set to their defaults. We chose not to use fallback to Trafilatura’s simple extraction baseline since it leaves most boilerplate intact, and we preferred sacrificing some documents but avoiding extra boilerplate in HPLT v2. Finally, we predict the language of the text using a modified version of the OpenLID model (Burchell et al., 2023; Burchell, 2024), where the Arabic dialects are combined under one macrolanguage label and the model training data has undergone improved pre-processing. These changes are intended to improve classification reliability. After text extraction, the dataset size reduces to 62 TB, 15 times smaller than the HTML data and 75 times smaller than the original web crawls.
@@ 4.2 Monolingual text processing
To respect robots.txt [^9] https://www.robotstxt.org/ rules specified by each domain, we use the extracted robots.txt records to identify patterns disallowing the crawlers we use.101010*, CCBot, ia-archiver We use the fst [^11] https://burntsushi.net/transducers/ tool to create a compressed index of URLs to exclude and use it to remove documents originating from these URLs.
@@ 5.1 Indirect quality indicators
We calculate descriptive statistics for HPLT v2 using the HPLT Analytics tool. [^17] https://github.com/hplt-project/data-analytics-tool We compare to the cleaned HPLT v1.2 dataset as the most comparable to our work.
@@ 5.2 Manual data inspection
To assess human-perceived quality, we manually inspected a random sample of documents from the cleaned monolingual datasets in 21 languages. Specifically, for each language spoken by the authors, we sampled 50 random documents extracted from each of the four groups of crawls: the older CC/IA crawls from 2012–2014, and the newer CC/IA crawls from 2017–2020. The main goal of this stratification was to compare the quality of texts we get depending on the crawl source and age, and select the most promising crawls for the next release of our datasets.
@@ 5.3 Register labels
We use the multilingual register classifier described in Henriksson et al. (2024) to label the entire monolingual HPLT v2. This classifier covers 16 languages and is based on an XLM-RoBERTa Large model Conneau et al. (2020a), fine-tuned on a multilingual web corpus manually annotated with register information. The classifier employs a hierarchical taxonomy with 25 register classes organized into 9 main categories (listed in Table 2).
@@
6 Corpora evaluation
In this section, we describe our empirical evaluation of the quality of the HPLT v2 datasets. We conduct this evaluation by employing the datasets as training material for several natural language processing (NLP) models.
@@ 6.1 Basic linguistic tasks and MLMs
We evaluate the trained MLMs on part-of-speech tagging, lemmatization and dependency parsing using the Universal Dependencies (UD) treebanks (de Marneffe et al., 2021), as well as named entity recognition (NER) using WikiAnn datasets (Pan et al., 2017). We compare to mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020b) models as multilingual baselines, and to HPLT v1.2 BERT models [^19] https://hf.co/collections/HPLT/hplt-bert-models-6625a8f3e0f8ed1c9a4fa96d as monolingual baselines. The performance is measured using the official CoNLL 2018 evaluation code (Zeman et al., 2018) for the UD tasks, and seqeval Nakayama (2018) balanced F1 score for the NER task.
@@ 6.2 NLU tasks and large generative LMs
Average results over the English benchmarks are presented in Figure 4. Our models trained on the HPLT v2 datasets reach similar performance to the models trained on FineWeb data and considerably outperform the models trained on HPLTv1.2. Specifically, the model trained on the cleaned subset of HPLT v2 is on par with the model trained on FineWeb data in downstream tasks, and shows improvement over the model trained on the deduplicated subset of HPLT v2. This implies that our cleaning approach has successfully improved the data quality with respect to these benchmarks.
@@ 6.3 Machine translation tasks
We evaluate the quality of the HPLT v2 parallel data by measuring the performance of MT models in two settings: as a complementary dataset to existing resources and as a stand-alone corpus.
Across all experiments, we carry out individual training for each language pair (to and from English) using the Transformer base architecture (Vaswani et al., 2017) and the Marian NMT toolkit (Junczys-Dowmunt et al., 2018). Data processing and training are streamlined with OpusPocus [^21] https://github.com/hplt-project/OpusPocus following the configuration of Arefyev et al. (2024). We evaluate all models on the FLORES-200 benchmark (NLLB Team et al., 2024) using BLEU (Papineni et al., 2002), chrF++ (Popović, 2017), and COMET-22-DA (Rei et al., 2022). We use sacrebleu’s implementation of the BLEU222222nrefs:1|case:mixed|eff:no|smooth:exp|version:2.5.1,
@@
and chrF++232323nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1 metrics (Post, 2018).
First, we show how HPLT v2 can be used as a complementary dataset together with existing collections. For this, we train models in three different data scenarios: 1) solely on the data from HPLT v2, 2) on the data from the Tatoeba Challenge (which includes most of the OPUS collection; Tiedemann, 2012, 2020), and 3) on a combination of the two.
Figure 6 summarises the average BLEU score for different settings, with detailed results in Tables 13 and 14 in the appendix. Our results show that models trained on HPLT v2 and Tatoeba perform on par on average. However, combining the two datasets results in a 7% relative increase in BLEU for both translation directions. This confirms that HPLT v2 offers non-overlapping content compared to other OPUS corpora, and as such is a valuable complementary resource for MT.
Figure 6: BLEU in different data scenarios for MT into and from English. We highlight the relative BLEU increase when adding HPLT v2 to Tatoeba.
Our second experiment investigates HPLT v2 as a stand-alone corpus. We compare the parallel portion of our dataset to HPLT v1.2 in order to see the effect of our improved data extraction pipeline discussed in Section 4. We consider the 10 languages which are covered in the parallel data of HPLT v2 and HPLT v1.2. Full results are given in the rightmost columns of Tables 13 and 14 in the appendix. Overall, The results show consistent improvements in BLEU scores for HPLT v2 across all models translating into English, with an average gain of 4.2 BLEU. For translations from English, the average gain is 3.5 BLEU, with 7 out of the 10 models showing better performance in both cases. These improvements clearly demonstrate the superior quality of HPLT v2, confirming its effectiveness as a resource for MT tasks.

### v3

@@ Abstract
Training state-of-the-art large language models requires vast amounts of clean and diverse textual data. However, building suitable multilingual datasets remains a challenge. In this work, we present HPLT v2, a collection of high-quality multilingual monolingual and parallel corpora, extending prior work of the HPLT project. The monolingual portion of the data contains 8T tokens covering 193 languages, while the parallel data contains 380M sentence pairs covering 51 languages. We document the entire data pipeline and release the code to reproduce it. We provide extensive analysis of the quality and characteristics of our data. Finally, we evaluate the performance of language models and machine translation systems trained on HPLT v2, demonstrating its value.
@@
High-Performance Language Technologies (HPLT)
Laurie Burchell1, Ona de Gibert2, Nikolay Arefyev3, Mikko Aulamo2, Marta Bañón4, Pinzhen Chen1, Mariia Fedorova3, Liane Guillou1, Barry Haddow1, Jan Hajič5, Jindřich Helcl5, Erik Henriksson6, Mateusz Klimaszewski1, Ville Komulainen6, Andrey Kutuzov3, Joona Kytöniemi6, Veronika Laippala6, Petter Mæhlum3, Bhavitvya Malik1, Farrokh Mehryary6, Vladislav Mikhailov3, Nikita Moghe7, Amanda Myntti6, Dayyán O’Brien1, Stephan Oepen3, Proyag Pal1, Jousia Piha6, Sampo Pyysalo6, Gema Ramírez-Sánchez4, David Samuel3, Pavel Stepachev1, Jörg Tiedemann2, Dušan Variš5, Tereza Vojtěchová5, Jaume Zaragoza-Bernabeu4 1University of Edinburgh, 2University of Helsinki, 3University of Oslo, 4Prompsit Language Engineering, 5Charles University, 6University of Turku, 7Amazonthanks: Work was done prior to joining Amazon. Contact: https://hplt-project.org
@@ 1 Introduction
In this paper, we introduce a new set of text corpora dubbed HPLT v2.111https://hplt-project.org/datasets/v2.0 It is extracted from 4.5 petabytes (PB) of the Internet Archive (IA) [^2] https://archive.org and Common Crawl (CC) [^3] https://commoncrawl.org data. We build on the work of de Gibert et al. (2024) (hereafter referred to as HPLT v1.2) with an improved extraction pipeline and a much larger set of input crawls to produce the HPLT v2 collection of monolingual and parallel corpora. To our knowledge, our new corpus is the only large-scale text collection extracted from the IA, apart from HPLT v1.2. We release HPLT v2 under the permissive Creative Commons Zero (CC0) license [^4] https://creativecommons.org/share-your-work/public-domain/cc0/. We do not claim ownership of any of the text from which this data has been extracted. and provide the code to replicate our pipeline. Our main contributions can be summarised as:
  • We release monolingual corpora covering 193 languages and containing approximately 52 trillion characters and 8 trillion tokens.
@@
Figure 1: The distribution of documents in the HPLT v2 cleaned dataset by language family and variety. Shortened ISO 639-3 language codes are used here; see the full plots at https://hplt-project.org/datasets/v2.0
@@ 2 Related work
The increasing data demands of state-of-the-art large language models have driven a rapid growth in both the number and the size of text corpora. We provide a summary of some well-known collections in Appendix A. Whilst LLMs trained on ostensibly English data have shown impressive multilingual capabilities Armengol-Estapé et al. (2022), of particular relevance to this work is the growing shift towards explicitly multilingual corpora. Compared with earlier efforts (e.g. OSCAR Suárez et al. (2019), CC-100 Conneau et al. (2020) and mC4 Xue et al. (2021)), more recent multilingual datasets cover increasing numbers of languages, e.g. CulturaX Nguyen et al. (2024a) and MADLAD-400 Kudugunta et al. (2024). HPLT v2 continues this trend by aiming for significant coverage of a wide range of languages. We note that the majority of previous multilingual datasets are sourced from CC, whereas much of HPLT v2 is composed of IA crawls. This means that HPLT v2 can be used in conjunction with these existing datasets as a complementary source.
@@
This work is a direct successor of our HPLT v1.2 dataset introduced in de Gibert et al. (2024). Compared to HPLT v1.2, we feature more data (21 billion vs. 5 billion documents) using an improved pipeline (Section 4), resulting in a significantly larger dataset (52 trillion characters compared to 42 trillion). The HPLT v2 collections are of higher quality than those in HPLT v1.2, as shown through comparative analysis and evaluation (Sections 5 and 6). For clarity, we list the main differences between HPLT v1.2 and HPLT v2 here:
  • The size of the source web collections is 2.5x larger, totalling 4.5 petabytes of compressed web data.
  • The text extraction pipeline uses Trafilatura rather than warc2text, which results in more efficient boilerplate removal.
  • Language identification uses a modified version of OpenLID rather than CLD2, increasing coverage from 75 to 193 language varieties.
  • Documents are annotated with their compliance to the robots.txt files of the corresponding websites, which can be used to filter out documents explicitly forbidden for crawling by website owners. The cleaned variant of the HPLT v2 dataset contains only robots.txt compliant documents.
  • Deduplication is done at collection level rather than globally.
  • Documents are annotated for PII information.
  • Document quality scores computed with web-docs-scorer rather than segment-level language model based scores
  • Simpler and more interpretable filtering and cleaning criteria.
@@ 3.1 Monolingual datasets
The monolingual portion of HPLT v2 covers 193 language varieties [^6] Language varieties are labelled with an ISO 639-3 code denoting the variety plus an ISO 15924 four-letter code denoting the script, separated by an underscore: e.g., gla_Latn. and is published in two variants: ‘deduplicated’ (21 terabytes) and ‘cleaned’ (15 terabytes). In the latter variant, the documents filtered by our cleaning heuristics (see Section 4.2) are excluded. For training LLMs, we recommend using the cleaned variant, but we also publish the datasets before cleaning (‘deduplicated’) so that it is possible to apply custom cleaning pipelines to the HPLT v2 data. In total, the deduplicated monolingual HPLT v2 datasets contain approximately 7.6 trillion white-space separated tokens and 52 trillion characters, extracted from 21 billion documents. HPLT v2 is published in the JSONL format, with one document per line.
@@ 3.2 Parallel datasets
We extract parallel data from the monolingual HPLT v2 to cover 50 languages paired with English. We aimed for a diverse range of language varieties and scripts in the low to medium resource range (listed in Table 5). We align these to English since this configuration has the highest potential for finding high-quality parallel data. We release our data in both XML and bitext format. [^7] https://opus.nlpl.eu/HPLT/corpus/version/HPLT
@@
DocHPLT v2
We also introduce context, by providing the documents that our bitext data comes from. These documents are annotated with both sentence and paragraph alignment. Overall, we provide 74,078,581 aligned documents covering all 51 languages. This ranges from 38,199,407 documents in English to 20,448 documents in Xhosa. [^8] https://opus.nlpl.eu/legacy/DocHPLT.php
@@
We leverage the English-centric HPLT v2 parallel resources to further create a multi-way parallel corpus, obtained by pivoting via English sentences. This corpus includes 1275 language pairs and contains over 16.7 billion parallel sentences. [^9] https://opus.nlpl.eu/MultiHPLT/corpus/version/MultiHPLT
A similar pivoting is also done for DocHPLT and we create the additional bitexts through sentences aligned to the same English document. However, anchoring sentences in aligned documents creates a different subset across the pivoted language pairs.
@@ 4.1 Text extraction from web crawls
This stage of the pipeline extracts the main textual content from HTML pages and groups it into language-specific subsets. It first parses the HTML pages into a tree representation. Next, it removes likely machine-translated texts by searching for indicative HTML tags and attributes. It then removes boilerplate (i.e. parts of a web page that do not contribute to its main content) using Trafilatura 1.8.0 (Barbaresi, 2021). Following hyperparameter experimentation, we set include_comments=False, include_tables=False, no_fallback=False and MIN_EXTRACTED_SIZE=0, with all other hyperparameters set to their defaults. We chose not to use fallback to Trafilatura’s simple extraction baseline since it leaves most boilerplate intact, and we preferred sacrificing some documents but avoiding extra boilerplate in HPLT v2. Finally, we predict the language of the text using a modified version of the OpenLID model [^13] https://data.statmt.org/lid/lid193_merged_arabics.bin (Burchell et al., 2023; Burchell, 2024), where the Arabic dialects are combined under one macrolanguage label and the model training data has undergone improved pre-processing. These changes are intended to improve classification reliability. After text extraction, the dataset size reduces to 62 TB, 15 times smaller than the HTML data and 75 times smaller than the original web crawls.
@@ 4.2 Monolingual text processing
To respect robots.txt [^14] https://www.robotstxt.org/ rules specified by each domain, we use the extracted robots.txt records to identify patterns disallowing the crawlers we use.151515*, CCBot, ia-archiver We use the fst [^16] https://burntsushi.net/transducers/ tool to create a compressed index of URLs to exclude and use it to remove documents originating from these URLs.
@@ 5.1 Indirect quality indicators
We calculate descriptive statistics for HPLT v2 using the HPLT Analytics tool. [^22] https://github.com/hplt-project/data-analytics-tool We compare to the cleaned HPLT v1.2 dataset (our previous release).
@@ 5.2 Manual data inspection
To assess human-perceived quality, we manually inspected a random sample of documents from the cleaned monolingual datasets in 22 languages. Specifically, for each language spoken by the authors, we sampled 50 random documents extracted from each of the four groups of crawls: the older CC/IA crawls from 2012–2014, and the newer CC/IA crawls from 2017–2020. The main goal of this stratification was to compare the quality of texts we get depending on the crawl source and age, and select the most promising crawls for the next release of our datasets.
@@ 5.3 Register labels
We use the multilingual register classifier described in Henriksson et al. (2024) to label the entire monolingual HPLT v2. This classifier covers 16 languages and is based on an XLM-R Large model (Conneau et al., 2020), fine-tuned on a multilingual web corpus manually annotated with register information. The classifier employs a hierarchical taxonomy with 25 register classes organized into 9 main categories (listed in Table 2). This label scheme is specifically designed for the linguistic characteristics of web texts and is therefore well-suited for our dataset.
@@
6 Empirical evaluation
In this section, we describe our empirical evaluation of the quality of the HPLT v2 monolingual corpora. We conduct this evaluation by employing the datasets as training material for several natural language processing (NLP) models.
@@ 6.1 Basic linguistic tasks and MLMs
We evaluate the trained MLMs on part-of-speech tagging, lemmatization and dependency parsing using the Universal Dependencies (UD) treebanks (de Marneffe et al., 2021), as well as named entity recognition (NER) using WikiAnn datasets (Pan et al., 2017). We compare to mBERT (Devlin et al., 2019) and XLM-R (Conneau et al., 2020) models as multilingual baselines, and to HPLT v1.2 BERT models [^24] https://hf.co/collections/HPLT/hplt-bert-models-6625a8f3e0f8ed1c9a4fa96d as monolingual baselines. The performance is measured using the official CoNLL 2018 evaluation code (Zeman et al., 2018) for the UD tasks, and seqeval Nakayama (2018) balanced F1 score for the NER task.
@@ 6.2 NLU tasks and large generative LMs
Average results over the English benchmarks are presented in Figure 4. Our models trained on the cleaned HPLT v2 datasets reach similar performance to the models trained on FineWeb data in downstream tasks, and considerably outperform the models trained on HPLT v1.2 and on the deduplicated subset of HPLT v2. This implies that our cleaning approach has successfully improved the data quality with respect to these benchmarks.
@@ 6.3 Machine translation tasks
We extrinsically evaluate the quality of the HPLT v2 parallel data by measuring the performance of MT models trained on it. This section considers two settings:
  • We compare HPLT v2 to its predecessor, HPLT v1.2, where we measure their performance in 10 overlapping languages.
  • We assess HPLT v2 as a complementary dataset to existing MT resources. We compare models in three data conditions: 1) solely on HPLT v2 parallel data; 2) on the data from Tatoeba Challenge (which includes most of the OPUS collection for each direction Tiedemann, 2012, 2020); and 3) on a combination of the two.
We carry out training for each individual language pair (to and from English) using the Transformer-base architecture (Vaswani et al., 2017) and the Marian NMT toolkit (Junczys-Dowmunt et al., 2018). Data processing and training are streamlined with OpusPocus [^26] https://github.com/hplt-project/OpusPocus following the configuration of Arefyev et al. (2024). We evaluate all models on the FLORES-200 benchmark (NLLB Team et al., 2024) using BLEU (Papineni et al., 2002), chrF++ (Popović, 2017), and COMET-22-DA (Rei et al., 2022). We use sacrebleu’s implementation of BLEU272727nrefs:1|case:mixed|eff:no|smooth:exp|version:2.5.1,
@@
and chrF++282828nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1 with signatures footnoted (Post, 2018). Full results for each metric and each direction are given in Table 14 and Table 15 in Appendix K.
xx-en en-xx BLEU COMET BLEU COMET HPLT v1.2 28.5 0.7943 24.4 0.7623 HPLT v2 32.7 0.8343 27.9 0.8137 HPLT v2 28.7 0.8144 23.4 0.7941 OPUS 29.6 0.8142 23.4 0.8074 HPLT v2+OPUS 30.5 0.8237 24.2 0.8083
Table 3: Average scores for the HPLT v1.2 and v2 comparison (top) and HPLT v2 as a complimentary resource to OPUS (bottom). Only numbers that are available for all models in a comparison are averaged.
We report average BLEU and COMET scores in Table 3. It is worth noting that we only average numbers for translation directions that are covered by all models in a comparison. The top half features a clear advantage of HPLT v2 over HPLT v1.2, reflecting the effect of our improved data extraction pipeline discussed in Section 4. Results in the bottom of Table 3 shows that HPLT v2 and OPUS perform comparably. However, combining the two datasets leads to improvements in both BLEU and COMET. This indicates that HPLT v2 contains non-overlapping content relative to existing OPUS corpora, making it a valuable complementary resource for MT.